OpenStudy (madmerc):

Simplify 49 -1/2 1/7 -1/7 7 -7

OpenStudy (madmerc):

@baru can you help with this one?

OpenStudy (anonymous):

Is it 49 * -1/2 ?

OpenStudy (madmerc):

yeah

OpenStudy (anonymous):

Okay, so all you have to do is multiply is straight across. \[\frac{ 49 }{ 1 }*\frac{ -1 }{ 2 }\]

OpenStudy (anonymous):

49 * -1 = -49 and 1 * 2 = 2 So your final answer is -49/2. In decimal form it would be -24.5

OpenStudy (madmerc):

wait. the equation is 49^-1/2

OpenStudy (madmerc):

\[49^{-1/2}\] like this

OpenStudy (anonymous):

Ohh, an exponent. Okay, sorry.

OpenStudy (madmerc):

that's okay I should've specified

OpenStudy (anonymous):

So we have to convert it to radical form it would be \[\frac{ 1 }{ \sqrt{49} }\] And that simplifies to 1/7 or 0.142 :)

OpenStudy (anonymous):

\[a ^{\frac{ x }{ n }}=\sqrt[n]{a ^{x}}\]The formula to convert to radical form is

OpenStudy (madmerc):

thank you!!